Find the maximum kinetic energy of the photoelectrons ejected when light of wavelength 350 nm is incident on a cesium surface. Work function of cesium = 1.9 eV.

Grade:10

1 Answers

Deepak Patra
9 years ago
Sol. λ = 350 nn = 350 * 10^-9 m ∅ = 1.9 eV Max KE of electrons =- =hC/λ - ∅ = 6.63 * 10^-34 *3 * 10^8/350 * 10^-9 * 1.6 * 10^-19 = 1.65 ev = 1.6 ev.